# Exploring a Patch-Wise Approach for Privacy-Preserving Fake ID Detection

### v1

@@ 3.2 Patch Generation
As described in Sec. 2, in this paper we explore fake ID detection based on patches, containing different information and sizes depending on the privacy restrictions. Regarding the pseudo- and fully-anonymized ID configurations, we covered the sensitive fields with pitch black rectangles (with the color code (0,0,0) in the RGB spectrum) using GNU Image Manipulator Program (GIMP) and EasyOCR [^5] https://github.com/JaidedAI/EasyOCR. After that, we used PyTorch’s unfold method to obtain the patches by specifying a stride with the same size as the patch. Patches with over 90% of its area with the (0,0,0) color in the RGB spectrum werediscarded. Additionally, the remaining patches were selected with a probability of $p=0.8$, to make reconstruction even more difficult. Examples of real and fake patches at different sizes can be seen in Fig. 3, where we can see that the patches alone do not contain any sensitive information.
In addition, as we plan to release the database, some mechanisms have been considered to avoid the reconstruction of the real IDs. First, only the pseudo- and fully-anonymized ID configurations are available for privacy reasons. Also, patches from all real and fake IDs are randomized in terms of the position and nomenclature. Finally, we would like to remark that the extracted patches from a single ID document are stored with a distinct code-name, so that this approach enables both single patch and full ID document evaluation.
@@ 4.1 Experimental Protocol
The experimental protocol carried out in the present study has been designed to analyze the feasibility of our proposed patch-wise approach for privacy-preserving fake ID detection.
@@
Patch Level ID Document Level
@@
Table 2: Performance in terms of EER (%) of our proposed patch-wise approach for privacy-preserving fake ID detection for the different patch size configurations and deep learning models. Evaluations at both patch and ID document levels are considered. For completeness, we also include in the “Full ID” column the results achieved for the case of introducing the whole ID picture to the deep learning models, instead of patches, as this is the most popular method in the literature.
@@
First, in Sec. 5.1 and Sec. 5.2, experiments are carried out considering our novel database. The purpose of this analysis is to evaluate our proposed method in terms of performance and privacy, comparing the results achieved with the traditional approach followed in the literature, i.e., introducing the whole ID document to the deep learning models.
@@
Regarding the data, we consider a balanced database composed of 30 real IDs and 30 fake IDs. Both types of PAIs (i.e., print and screen) are used for training our proposed fake ID detection method. Table 1 provides a summary of the total number of IDs (shown as “Full”) and patches available depending on the anonymization configuration and the size of the patches. As can be seen, it is important to highlight that the amount of training data changes based on the selected privacy configuration, i.e., anonymization and patch size. Therefore, these experiments will evaluate the influence of anonymization and patch size configurations in the final trade-off of performance and privacy. Regarding the experimental protocol, our database is divided into development (80% of the real/fake IDs) and final evaluation (remaining 20% of the real/fake IDs) datasets. This considers a realistic scenario with unseen IDs for testing and ID templates (first version of the Spanish ID is only seen in the final evaluation, not training). Evaluation is performed at both patch and full ID document level.
@@
After this first analysis using only our novel database, we evaluate in Sec. 5.4 the performance of our proposed patch-wise approach for privacy-preserving fake ID detection under a cross-database scenario. This scenario intends to analyze the generalization ability of the proposed fake ID detection method to unseen PAIs and databases not considered for training, which is expected to be the typical scenario in real applications. In particular, the optimal configuration of our proposed method, discussed in Sec. 5.3 and trained only on our novel database, is evaluated using PAIs from a different public database, DLC-2021 [19]. This database includes physical PAIs such as print and screen, which are created following similar procedures as ours, but captured in different conditions. For example, in DLC-2021, the acquisition is performed using different devices (i.e. Samsung S10 and iPhone XR vs. Redmi 9C NFC), distance, and light conditions, among others. More details and differences between both databases are included in Sec. 5.4. Concretely, from DLC-2021 we consider 1,500 fake document samples (balanced among the different PAI species), which are processed considering the best patch configuration (see Sec. 5.3). The real patches are extracted from our novel database (i.e., the evaluation dataset) as there are no public databases that contain real ID samples.
@@ 4.2 Evaluation Metrics
Similar to previous approaches presented in the literature [23], we consider the ISO/IEC 30107-3 standard [^6] https://www.iso.org/standard/79520.html for the evaluation of fake ID detection technology: the Bona-fide Presentation Classification Error Rate (BPCER) and the Attack Presentation Classification Error Rate (APCER). The BPCER metric tells how many bona-fide samples (i.e., real IDs) are incorrectly classified as attacks (i.e., fake IDs). The APCER metric represents the same thing as the BPCER but for the attacks (i.e., fake IDs) that are incorrectly classified as bona fide (i.e., real IDs). In addition, the Equal Error Rate (EER) metric, which gives the error rate at the operational point $\tau$ where the ${\leavevmode BPCER}(\tau)$ and ${\leavevmode APCER}(\tau)$ metrics are equal, is another popular metric considered in the literature to compare performance. Finally, as described in Sec. 2, our proposed method is designed to detect real/fake patches as well as the whole ID document. This is referred as “Patch Level” or “ID Document Level”.
@@ 5 Experimental Results
This section explores the feasibility of our proposed patch-wise approach for privacy-preserving fake ID detection. In particular, in Sec. 5.1 we analyze the performance of our proposed method in terms of the patch size (i.e., $128\times 128$, $64\times 64$, and $32\times 32$). Sec. 5.2 evaluates the performance of the proposed method in terms of the anonymization setup (i.e., non-anonymized ID, pseudo-anonymized ID, and fully-anonymized ID). Then, in Sec. 5.3, we select the optimal configuration of our proposed method in terms of performance and privacy. Finally, in Sec. 5.4 we evaluate the performance of our optimal fake ID detection method under a cross-database scenario. This scenario intends to analyze the generalization ability of the proposed fake ID detection method to unseen PAIs and databases not considered for training. Evaluations at both patch and ID document level are considered in the analysis.
@@
Table 2 shows the performance in terms of EER (%) of our proposed method for the different patch size configurations and deep learning models. Evaluations at both patch and ID document level are considered. For completeness, we also include in the “Full ID” column the results achieved for the case of introducing the whole ID picture to the models, instead of patches, as this is the most popular method in the literature. In this first analysis, we consider the case of having all information of the ID document available (i.e., non-anomymized ID configuration).
Patch Level ID Document Level
@@
ResNet-18 ($128\times 128$) 16.62 16.80 11.83 0.00 0.00 0.00
ViT-B/16 ($128\times 128$) 18.61 19.72 20.97 0.00 0.00 0.00
DINOv2 ($128\times 128$) 13.81 17.76 15.59 0.00 0.00 0.00
Avg. EER ($128\times 128$) 16.34 18.09 16.13 0.00 0.00 0.00
ResNet-18 ($64\times 64$) 20.88 22.44 22.60 0.00 0.00 0.00
ViT-B/16 ($64\times 64$) 19.31 20.08 21.51 0.00 0.00 0.00
DINOv2 ($64\times 64$) 14.98 15.38 14.83 0.00 0.00 0.00
Avg. EER ($64\times 64$) 18.39 19.3 19.64 0.00 0.00 0.00
@@
Table 3: Performance in terms of EER (%) of our proposed method for the different anonymization configurations (i.e., non-anonymized ID, pseudo-anonymized ID and fully-anonymized ID), the two best patch size configurations ($128\times 128$ and $64\times 64$), and all deep learning models. Evaluations at both patch and ID document level are considered.
@@
Analyzing the performance of our proposed method at patch level, we can see that, in general, the best patch size configuration is $128\times 128$, achieving an average EER of 16.34%. The EER increases as the patch size is reduced, with values of 18.39% and 25.27% EER for the $64\times 64$ and $32\times 32$ configurations, respectively. These results are also interesting from the point of view of the amount of training data. For example, for the $128\times 128$ configuration, in this experiment we have 9,520 patches in total whereas for the $32\times 32$ configuration this value increases to 144,160 patches. The results achieved in this particular experiment reject the hypothesis that more patches of smaller size may perform better than fewer patches of bigger size.
@@
Also, it is important to highlight the different performance of the deep learning models depending on the patch size configuration. DINOv2 achieves the best performance with 13.81% EER for the $128\times 128$ configuration, an absolute improvement of 2.79% and 4.80% EER in comparison to the ResNet-18 and ViT-B/16 models, respectively. This performance improvement may be produced due to the objective of the training procedure of DINOv2 is to match the embeddings distribution from images with full context, to local patches of said images. By following this procedure, DINOv2 is able to align both local and global features from a single image, and potentially benefits from local information given as patches in our proposed method.
@@
It is also interesting to analyze the optimal trade-off between performance and privacy, as the patch size contributes directly to the privacy as much or less information of the subject would be visible in the patch images, as can be seen in Fig. 3. In this sense, it is surprising to see that DINOv2 performs quite similar for the $64\times 64$ configuration, with an EER absolute increasing of just 1.17% (13.81% vs. 14.98%). A similar trend is observed for the ViT-B/16 model, where its difference in performance between the patch size configurations is just 0.7% EER (18.61% vs. 19.31%). These results seem to suggest that deep learning models based on ViT architectures provide features that are more robust to patch size. In the case of ResNet-18, it achieves a good performance for the $128\times 128$ configuration, but lacks of robustness when the patch size decreases, e.g., from 16.62% to 30.17% EERs for the $128\times 128$ and $32\times 32$ configurations, respectively.
@@
Finally, we include in Table 2 the performance of our proposed method at the whole ID document level. For completeness, we also include in the “Full ID” column the results achieved for the case of introducing the whole ID picture to the deep learning models, instead of patches, as this is the most popular method in the literature. In general, very good results can be achieved using our proposed method, with 0% EER results in most configurations and deep learning models. These results are much better compared to the “Full ID” traditional case considered in the state of the art, e.g., for the case of DINOv2 the EER increases to 33.3% for the “Full ID” scenario. These results confirm the feasibility of our proposed patch-wise approach for privacy-preserving fake ID detection.
@@
In addition to the different patch size configurations studied before, the present section analyzes the effect of the different anonymization configurations in the final performance. Similar to the previous analysis, the aim is to obtain a trade-off between performance and privacy.
Table 3 shows the performance in terms of EER (%) of our proposed method for the different anonymization configurations (i.e., non-anonymized ID, pseudo-anonymized ID, and fully-anonymized ID), and the two best patch size configurations ($128\times 128$ and $64\times 64$), and all deep learning models. Evaluations at both patch and ID document level are considered.
Analyzing the different anonymization configurations, in general we can observe in all deep learning models an EER increasing from the non-anonymized to the fully-anonymized configurations, e.g., for the $64\times 64$ patch size, the EER increases on average from 18.39% to 19.64% EER. These results make sense as more patches and with additional fake patterns in the sensitive information are available for the non-anonymized configuration. Nevertheless, the performance is in general similar for all anonymization configurations, being possible to reduce the amount of sensitive information to detect if an ID document is real or fake, e.g., for the DINOv2 model and $64\times 64$ patch size, EER values of 14.98% and 14.83% are achieved for the non- and fully-anonymized configurations, respectively.
Finally, for completeness, we also include in Table 3 the performance of our proposed method at the whole ID document level. Again, very good results can be achieved using our proposed method in all anonymization configurations, with 0% EER results in both $128\times 128$ and $64\times 64$ patch size configurations and all deep learning models. These results confirm again the feasibility of our proposed patch-wise approach for privacy-preserving fake ID detection.
@@ 5.3 Proposed Method: Optimal Configuration
The results achieved in Sec. 5.1 and Sec. 5.2 have proved the feasibility of our proposed method in comparison to the traditional one, i.e., introducing the whole ID picture to the deep learning model. Several configurations have been studied in terms of patch size and anonymization with the purpose of selecting an optimal trade-off between performance and privacy, due to the sensitive information included in ID documents.
Taking this into account, we have decided to select as our optimal configuration the DINOv2 model with the patch size $64\times 64$ and the fully-anonymized ID setup. This is a privacy-preserving configuration as no sensitive information is considered in the analysis. In addition, as can be seen in Fig. 3, images with a patch size of $64\times 64$ reveal far less information of the subject than the $128\times 128$ configuration. For this particular configuration, DINOv2 achieves EER values of 14.83% at patch level and 0% at ID document level.
Figure 4: Examples of fake patches from the DLC-2021 [19].
@@ 5.4 Cross-Database Scenario
This section evaluates the generalization ability of the proposed fake ID detection method to unseen PAIs and databases not considered for training, which is expected to be the typical scenario in real applications. Concretely, we consider the optimal configuration described in Sec. 5.3, trained only with our novel database. For the final evaluation, we consider fake ID documents from a different database, DLC-2021 [19]. This database includes physical PAIs such as print (glossy, color, and gray) and screen. Graphical examples of $64\times 64$ fake patches extracted from DLC-2021 are included in Fig. 4. Table 4 shows the performance in terms of EER (%) of DINOv2 ($64\times 64$, fully-anonymized ID) for this database.
Before analyzing the results, we would like to remark the difficulty of the scenario as: i) different types of physical PAIs are considered in the evaluation of the fake ID detection method, ii) different types of templates and documents are considered in the analysis (ID card and passport), and from different countries such as Albania, Findland, Estonia, etc., and iii) different types of acquisitions are considered in terms of smartphones (iPhone XR y Samsung S10), resolution, distance of the camera, angles, etc. Additionally, we would like to remark that the DLC-2021 database contains Spain ID documents, but their templates belong to the first version, which is not seen while training our proposed method, as commented in the experimental protocol, see Sec. 4.1.
PAI Class Patch Level ID Document Level
@@
Table 4: Performance in terms of EER (%) of DINOv2 ($64\times 64$, fully-anonymized ID) for the DLC-2021 [19]. This database is not considered for training our proposed method. Fake IDs from different PAIs are considered in the analysis.
As can be seen in Table 4, our proposed method is able to generalize well to patches extracted from different PAIs. In particular, the proposed method achieves on average 13.91% EER for the analysis at patch level, similar to the performance achieved in our database (i.e., 14.83% EER). This suggests that when physical PAIs are used, our method remains robust, effectively detecting fake patches from different distributions. Finally, if we analyze the performance at the whole ID document level, we can see that our proposed method is able to detect real/fake IDs without mistakes (0% EER), proving a good generalization ability.

### v2

@@ 3.2 Patch Generation
As described in Sec. 2, in this paper we explore the detection of fake IDs based on patches, containing different information and sizes depending on privacy restrictions. Regarding the pseudo- and fully-anonymized ID configurations, we covered the sensitive fields with pitch black rectangles (with the color code (0,0,0) in the RGB spectrum) using the GNU Image Manipulator Program (GIMP) and EasyOCR [^5] https://github.com/JaidedAI/EasyOCR. After that, we used PyTorch’s unfold method to obtain the patches by specifying a stride with the same size as the patch. Patches with more than 90% of its area with the (0,0,0) color in the RGB spectrum were discarded. Furthermore, the remaining patches were selected with a probability of $p=0.8$, making reconstruction even more difficult. Examples of real and fake patches of different sizes can be seen in Fig. 3, where we can see that the patches alone do not contain sensitive information.
In addition, as we plan to release the database, some mechanisms have been considered to avoid the reconstruction of the real IDs. First, only the pseudo- and fully-anonymized ID configurations are available for privacy reasons. In addition, patches from all real and fake IDs are randomized in terms of position and nomenclature. Finally, we remark that the extracted patches from a single ID are stored with a distinct code name, so that this approach enables both single patch and full ID evaluation.
@@ 4.1 Experimental Protocol
The experimental protocol has been designed to analyze the feasibility of our proposed FakeIDet method.
@@
Patch Level ID Level
@@
Table 2: Performance in terms of EER (%) of FakeIDet for the different patch size configurations and deep learning models. Evaluations at both patch and ID levels are considered. For completeness, we also include in the “Full ID” column the results achieved for the case of introducing the whole ID picture to the deep learning models, instead of patches, as this is the most popular method in the literature.
@@
First, in Sec. 5.1 and Sec. 5.2, experiments are carried out considering our novel database. The purpose of this analysis is to evaluate our proposed method in terms of performance and privacy, comparing the results achieved with the traditional approach followed in the literature, i.e., introducing the whole ID to the deep learning models.
@@
Regarding the data, we consider a balanced database composed of 30 real IDs and 30 fake IDs. Both types of PAIs (i.e., print and screen) are used to train our proposed fake ID detection method. Table 1 provides a summary of the total number of IDs (shown as “Full”) and patches available depending on the anonymization configuration and the size of the patches. As can be seen, it is important to highlight that the amount of training data changes based on the selected privacy configuration, i.e., anonymization and patch size. Therefore, these experiments will evaluate the influence of anonymization and patch size configurations in the final trade-off of performance and privacy. Regarding the experimental protocol, our database is divided into the development (80% of the real/fake IDs) and final evaluation (remaining 20% of the real/fake IDs) dataset. This considers a realistic scenario with unseen IDs for testing (the first version of the Spanish ID is only seen in the final evaluation, not in training). Evaluation is performed at both patch and full ID levels.
@@
After this first analysis using only our novel database, we evaluate in Sec. 5.4 the performance of our proposed patch-wise approach for privacy-preserving fake ID detection under a cross-database scenario. This scenario aims to analyze the generalizability of FakeIDet to unseen databases, which is expected to be the typical scenario in real applications. In particular, the optimal configuration of our proposed method, discussed in Sec. 5.3 and trained only on our new database is evaluated using PAIs from a different public database, DLC-2021 [25]. This database includes physical PAIs such as print and screen, which are created following similar procedures as ours but captured under different conditions. For example, in DLC-2021, the acquisition is performed using different devices (i.e. Samsung S10 and iPhone XR vs. Redmi 9C NFC), distance, and light conditions, among others. More details and differences between both databases are included in Sec. 5.4. Concretely, from DLC-2021 we consider 1,500 fake document samples (balanced among the different PAI species), which are processed considering the best patch configuration (see Sec. 5.3). The real patches are extracted from our novel database (i.e., the evaluation dataset) as there are no public databases that contain real ID samples.
@@ 4.2 Evaluation Metrics
[…]
@@ 5 Experimental Results
We explore the feasibility of our proposed patch-wise approach for privacy-preserving fake ID detection. In particular, in Sec. 5.1 we analyze the performance of FakeIDet in terms of patch size (i.e., $\times 128$ 128, $\times 64$ 64, and $\times 32$ 32). Sec. 5.2 evaluates the performance of FakeIDet in terms of the anonymization setup (i.e., non-, pseudo-, and fully-anonymized ID). Then, in Sec. 5.3, we select the optimal configuration of FakeIDet in terms of performance and privacy. Finally, in Sec. 5.4 we evaluate the performance of our optimal FakeIDet in a cross-database scenario. This scenario aims to analyze the generalizability of the proposed fake ID detection method to unseen PAIs and databases not considered for training. Evaluations at both the patch level and the identification document level are considered in the analysis.
@@
Table 2 shows the performance in terms of EER (%) of our FakeIDet method for the different patch size configurations and deep learning models. Evaluations are performed at both the patch level and the ID level. For completeness, we also include in the “Full ID” column the results achieved for the case of introducing the whole ID picture to the models, instead of patches, as this is the most popular method in the literature. In this first analysis, we consider the case of having all information of the ID available (i.e., non-anomymized ID configuration).
Patch Level ID Level
@@
ResNet-18 ($\times 128$ 128) 16.62 16.80 11.83 0.00 0.00 0.00
ViT-B/16 ($\times 128$ 128) 18.61 19.72 20.97 0.00 0.00 0.00
DINOv2 ($\times 128$ 128) 13.81 17.76 15.59 0.00 0.00 0.00
Avg. EER ($\times 128$ 128) 16.34 18.09 16.13 0.00 0.00 0.00
ResNet-18 ($\times 64$ 64) 20.88 22.44 22.60 0.00 0.00 0.00
ViT-B/16 ($\times 64$ 64) 19.31 20.08 21.51 0.00 0.00 0.00
DINOv2 ($\times 64$ 64) 14.98 15.38 14.83 0.00 0.00 0.00
Avg. EER ($\times 64$ 64) 18.39 19.3 19.64 0.00 0.00 0.00
@@
Table 3: Performance in terms of EER (%) of our FakeIDet method for the different anonymization configurations (i.e., non-, pseudo-, and fully-anonymized ID), the two best patch size configurations ($\times 128$ 128 and $\times 64$ 64), and all deep learning models. Evaluations at both patch and ID level are considered.
@@
Analyzing the performance of our proposed method at the patch level, we can see that, in general, the best patch size configuration is $\times 128$ 128, achieving an average EER of 16.34%. The EER increases as the patch size is reduced, with values of 18.39% and 25.27% EER for the $\times 64$ 64 and $\times 32$ 32 configurations, respectively. These results are also interesting from the point of view of the amount of training data. For example, for the $\times 128$ 128 configuration, in this experiment we have 9,520 patches in total, whereas for the $\times 32$ 32 configuration this value increases to 144,160 patches. The results achieved in this particular experiment reject the hypothesis that more patches of smaller size may perform better than fewer patches of larger size.
@@
Also, it is important to highlight the different performance of deep learning models depending on the patch size configuration. DINOv2 achieves the best performance with 13.81% EER for the $\times 128$ 128 configuration, an absolute improvement of 2.79% and 4.80% EER compared to the ResNet-18 and ViT-B/16 models, respectively. This performance improvement may be produced due to the objective of the DINOv2 training procedure being to match the embedding distribution from images with full context to local patches. By following this procedure, DINOv2 is able to align both local and global features from a single image, and potentially benefits from local information given as patches in our proposed method.
@@
It is also interesting to analyze the optimal trade-off between performance and privacy, as patch size directly contributes to privacy, since much or less information of the subject would be visible in patch images, as can be seen in Fig. 3. In this sense, it is surprising to see that DINOv2 performs quite similar for the $\times 64$ 64 configuration, with an absolute increase in EER of just 1.17% (13.81% vs 14.98%). A similar trend is observed for the ViT-B/16 model, where its difference in performance between patch size configurations is only 0.7% EER (18.61% vs. 19.31%). These results suggest that deep learning models based on ViT architectures provide features that are more robust to patch size. In the case of ResNet-18, it achieves good performance for the $\times 128$ 128 configuration, but lacks robustness when the patch size decreases, e.g., from 16.62% to 30.17% EER for the $\times 128$ 128 and $\times 32$ 32 configurations, respectively.
@@
Finally, in Table 2 we include the performance of the FakeIDet method at the whole ID level. For completeness, we also include in the “Full ID” column the results achieved for the case of introducing the whole ID picture to the deep learning models, instead of patches, as this is the most popular method in the literature. In general, very good results can be achieved using our proposed method, with 0% EER results in most configurations and deep learning models. These results are much better compared to the “Full ID” traditional case considered in the state of the art, e.g., for the case of DINOv2 the EER increases to 33.3% for the “Full ID” scenario. These results confirm the feasibility of our proposed patch-wise approach for privacy-preserving fake ID detection.
@@
In addition to the different patch size configurations studied before, the present section analyzes the effect of the different anonymization configurations in the final performance. Similarly to the previous analysis, the aim is to obtain a trade-off between performance and privacy.
Table 3 shows the performance in terms of EER (%) of our FakeIDet method for the different anonymization configurations (i.e., non-, pseudo-, and fully-anonymized ID), and the two best patch size configurations ($\times 128$ 128 and $\times 64$ 64), and all deep learning models. Evaluations are considered at both the patch and the ID levels.
Analyzing the different anonymization configurations, in general we can observe in all deep learning models an EER increasing from the non-anonymized to the fully-anonymized configurations, e.g., for the $\times 64$ 64 patch size, the EER increases on average from 18.39% to 19.64% EER. These results make sense as more patches and with additional fake patterns in the sensitive information are available for the non-anonymized configuration. Nevertheless, the performance is in general similar for all anonymization configurations, being possible to reduce the amount of sensitive information to detect if an ID is real or fake, e.g., for the DINOv2 model and $\times 64$ 64 patch size, EER values of 14.98% and 14.83% are achieved for the non- and fully-anonymized configurations, respectively.
Finally, for completeness, we also include in Table 3 the performance of our proposed method at the whole ID level. Again, very good results can be achieved using our proposed method in all anonymization configurations, with 0% EER results in both $\times 128$ 128 and $\times 64$ 64 patch size configurations and all deep learning models. These results confirm again the feasibility of our proposed patch-wise approach for privacy-preserving fake ID detection.
@@ 5.3 Proposed Method: Optimal Configuration
The results achieved in Sec. 5.1 and Sec. 5.2 have proved the feasibility of our FakeIDet method compared to the traditional one, that is, introducing the whole ID picture into the deep learning model. Several configurations have been studied in terms of patch size and anonymization with the purpose of selecting an optimal trade-off between performance and privacy due to the sensitive information included in IDs.
Taking this into account, we have decided to select as our optimal configuration the DINOv2 model with the patch size $\times 64$ 64 and the fully-anonymized ID setup. This is a privacy-preserving configuration, as sensitive information is not considered in the analysis. In addition, as can be seen in Fig. 3, images with a patch size of $\times 64$ 64 reveal much less information on the subject than the $\times 128$ 128 configuration. For this particular configuration, DINOv2 achieves EER values of 14.83% at the patch level and 0% at the ID level.
Figure 4: Examples of fake patches from DLC-2021 [25].
@@ 5.4 Cross-Database Scenario
This section evaluates the generalizability of our proposed FakeIDet method to unseen PAIs and databases not considered for training, which is expected to be the typical scenario in real applications. Concretely, we consider the optimal configuration described in Sec. 5.3, trained only with our novel database. For the final evaluation, we consider fake IDs from a different database, DLC-2021 [25]. This database includes physical PAIs such as print (glossy, color, and gray) and screen. Graphical examples of $\times 64$ 64 fake patches extracted from DLC-2021 are included in Fig. 4. Table 4 shows the performance in terms of EER (%) of DINOv2 ($\times 64$ 64, fully-anonymized ID) for this database.
Before analyzing the results, we would like to remark the difficulty of the scenario as: i) different types of physical PAIs are considered in the evaluation of the fake ID detection method, ii) different types of templates and documents are considered in the analysis (ID card and passport), and from different countries such as Albania, Findland, Estonia, etc., and iii) different types of acquisitions are considered in terms of smartphones (iPhone XR y Samsung S10), resolution, distance of the camera, angles, etc. Additionally, we would like to remark that the DLC-2021 database contains Spain IDs, but their templates belong to the first version, which is not seen while training our proposed method, as commented in the experimental protocol, see Sec. 4.1.
PAI Class Patch Level ID Level
@@
Table 4: Performance in terms of EER (%) of our FakeIDet method ($\times 64$ 64, fully-anonymized ID) for the DLC-2021 [25] database, which is not considered for training our proposed method. Fake IDs from different PAIs are considered in the analysis.
As can be seen in Table 4, our proposed FakeIDet method is able to generalize well to patches extracted from different PAIs. In particular, the proposed method achieves on average 13.91% EER for the analysis at patch level, similar to the performance achieved in our database (i.e., 14.83% EER). This suggests that when physical PAIs are used, our method remains robust, effectively detecting fake patches from different distributions. Finally, if we analyze the performance at the whole ID level, we can see that our proposed method is able to detect real/fake IDs without mistakes (0% EER), proving a good generalization ability.